atives are mutually exclusive and have infinite useful lives, determine shict of te llowing alternative should be selected for the investment using internal rate of re method (MARR is given as 10%) (20 points) A Project B Project C Projet Investment 10.000.000 8500000 7.600.00 Cost TL TL TL Salvage value eenues | 950.000 TL | 900.000 TLI800000L 5.000.0001.000000TL TL TL
Answer #1

Salvage value is not relevant when useful life is infinite

Project A:

Equating PV = discounted value of annual revenues at IRR

10000000 = 950000/IRR

IRR = 9.5%

Project B:

8500000 = 900000/IRR

IRR = 10.59%

Project C:

7600000 = 800000/IRR

IRR = 10.53%

Project B should be selected owing to highest IRR > MARR.